How do I solve?
7-2(3x - 1)=1 - 3(x-1)
Thanks

Answer by ptaylor(2198)

How do I solve?
7-2(3x - 1)=1 - 3(x-1)
Thanks

Step 1 ----get rid of parens by multiplying out and we get:
7-6x+2=1-3x+3
Step 2----simplify by collecting like terms on each side of the equation and we have:
9-6x=4-3x
Step 3----we will add 3x to both sides and, at the same time, we will subtract 9 from both sides. This will put the numbers on one side and the unknown on the other and that's what we want.
9-9-6x+3x=4-9-3x+3x
Step 4---we will collect like terms again:
-3x=-5
Step 5----divide both sides by -3
x=5/3---------answer
(Note: the above steps can vary depending on the equation being solved)
CK
7-2(3x - 1)=1 - 3(x-1)
substitute 5/3 for x
7-2((15/3)-1)=1-3((5/3)-1)=
7-30/3+2=1-15/3+3=
7-10+2=1-5+3
-1=-1
Hope this helps-----ptaylor